# Towards a logic of affordances

### v1

@@ 5. Reasoning with affordances: modal-style operators
Having conceptualized affordances, we then turn to and focus on the problem of reasoning with them as ternary relations. As we will see, the equivalence classes of entities from three different sorts will play a crucial role in incorporating the granularity of perceptions of the world into the formal framework. In this section however, we begin with the definitions of crisp modal-style operators. As such these are independent from the affordance environment above, and can be laid out in a completely abstract setting. To this end, suppose that
@@
Their dual operators are, respectively, $\varphi]_{E}$ and $\left\langle\left\langle\varphi\right\rangle\right\rangle_{E}$ , where
@@
Let us use one more time the interpretation from Example 5.1. Then:
@@
The second point confirms the awkwardness of the dual of the sufficiency operator.
Analogously to [9, Lemma 5.2] we have the wMIA conditions for the operators associated with respective sets.
@@
Let us look at the concrete example of a ternary relation in the form of betweenness that was studied by us in the setting of possibility-sufficiency algebras in [9]. $B(a,o,e)$ is interpreted as: the point $o$ is between (in the non-strict sense) the points $a$ and $e$. Then $e\in\left\langle B\right\rangle\rangle(X,Y)$ if and only if either there are $x\in X$ and $y\in Y$ such that $B(x,e,y)$ or either there are $x\in X$ and $y\in Y$ such that $\neg B(x,e,y)$. Then $\left\langle B\right\rangle\rangle(X,Y)=\emptyset$ if and only if at least one of the sets is empty, and $\left\langle B\right\rangle\rangle(X,Y)$ is the set of all points otherwise. $\dashv$
@@
1. Let $e\in E$. It is obvious that for any $x\in X$ and $o\mathtt{ON}O$, if $\left\langle x,o,e\right\rangle\in\varphi$, then $x\in X$ or $o\in O$. From (5.7), $e\in[\varphi]_{E}(X,O)$, i.e. $\varphi]_{E}(X,O)=E$. In an analogous way we obtain that $\varphi]_{E}(A,Y)=E$.
@@ 6. The approximation operators
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(a)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\overline{\varphi}\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;([a]\times[o]\times[e])\cap\varphi\neq\emptyset.$
@@
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(a)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\underline{\varphi}\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;[a]\times[o]\times[e]\subseteq\varphi.$
@@
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(a)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\overline{\varphi}$
@@
A rough affordance is a pair $\left\langle\underline{\varphi},\overline{\varphi}\right\rangle$ where $\varphi$ is an affordance relation.
@@
6.1. Upper approximation operators
With an affordance $\varphi$ we can associate its upper approximation $\overline{\alpha}_{S}(\varphi)$ on every coordinate $S\in\{A,O,E\}$ (i.e., we have approximations for, respectively, actors, objects, end environments):
@@ 6.1. Upper approximation operators
$\overline{\alpha}(\varphi)_{S}=\overline{\alpha}\left(\overline{\varphi}\right)_{S}.$
@@
($\subseteq$) Let $e\in\overline{\alpha}(\varphi)_{E}(X,Y)$, i.e., $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\varphi\neq\emptyset$. Then, by Lemma 6.2 it must be the case that $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. So $e\in\overline{\alpha}(\overline{\varphi})_{E}(X,Y)$.
($\supseteq$) For the other direction, assume that $e\in\overline{\alpha}(\overline{\varphi})_{E}(X,Y)$, that is $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. Let $x\in\overline{X}$, $y\in\overline{Y}$, and $z\in[e]$ be such that $([x]\times[y]\times[z])\cap\varphi\neq\emptyset$. Let $\left\langle a,o,u\right\rangle$ be a triple which is in the intersection. But then $[a]=[x]$ and $[o]=[y]$ and $[u]=[e]$, and so $a\in\overline{X}$, $o\in\overline{Y}$, and in consequence $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. Thus $e\in\overline{\alpha}(\varphi)_{E}(X,Y)$, as desired. ∎
@@
  • $r\in\overline{\alpha}^{1}_{A}(\varphi)(S,P)$ means that a dog similar to Robin fetches a stick in a dog park.
  • $r\in\overline{\alpha}^{2}_{A}(\varphi)(S,P)$ means that Robin fetches a stick-like object in a dog park.
  • $r\in\overline{\alpha}^{3}_{A}(\varphi)(S,P)$ means that Robin fetches a stick in an environment resembling a dog park.
  • $r\in\overline{\alpha}^{12}_{A}(\varphi)(S,P)$ means that in a dog park a dog similar to Robin fetches an object similar to a stick.
  • $r\in\overline{\alpha}^{13}_{A}(\varphi)(S,P)$ means that in an environment resembling a dog park a dog similar to Robin fetches a stick.
  • $r\in\overline{\alpha}^{23}_{A}(\varphi)(S,P)$ means that Robin fetches something similar to a stick in an environment resembling a dog park.
Thus, playing around with the coordinates we can express roughness of various aspects of the scenarios we put into focus. $\dashv$
For any $i\in\{1,2,3,12,13,23\}$ by Lemma 6.2 we have that
@@
However, the reverse inclusions fail in general, as we show below.
@@
We stick to the interpretation from the previous two examples, and we consider the case in which $i\coloneqq 23$. Then, the statement $r\in\overline{\alpha}^{23}_{A}(\overline{\varphi})(S,P)$, i.e., $(\{r\}\times\overline{S}\times\overline{P})\cap\overline{\varphi}\neq\emptyset$, according to the definition of a rough affordance says that there is a triple $\left\langle r,s,p\right\rangle\in\{r\}\times\overline{S}\times\overline{P}$ such that $([r]\times[s]\times[p])\cap\varphi\neq\emptyset$. So it means that there is a triple of entities $\left\langle x,y,z\right\rangle$ such that $x$ is a pet similar to Robin, $y$ is similar to $s$ which resembles a stick, and $z$ is similar to $p$ which resembles a dog park. Since in this case, similarity is an equivalence relation we have that $y$ resembles a stick and $z$ resembles a dog park. In consequence, $r\in\overline{\alpha}^{23}_{A}(\overline{\varphi})(S,P)$ means that a dog resembling Robin is fetching a stick-like object in a dog-park-like environment. So we can conclude that $r\in\overline{\alpha}_{A}(\varphi)(S,P)$, but not that $r\in\overline{\alpha}_{A}^{23}(\varphi)(S,P)$, that is we cannot obtain certainty with respect to the pet carrying out the action. $\dashv$
@@ 6.2. Lower approximation operators
For each $S\in\{A,O,E\}$ we have that
$\underline{\alpha}(\varphi)_{S}(X,Y)=\underline{\alpha}(\underline{\varphi})_{S}(X,Y).$
@@
This time, the inclusion $\supseteq$ is immediate in light of Lemma 6.2. For $\subseteq$, assume that $e\in\underline{\alpha}(\varphi)_{E}(X,Y)$, i.e., $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$. Let $\left\langle a,o,f\right\rangle$ be from $\underline{X}\times\underline{Y}\times[e]$. Thus $a]\subseteq X$, $o]\subseteq Y$, and $[f]=[e]$. So, $a]\subseteq\underline{X}$, $o]\subseteq\underline{Y}$, and $a]\times[o]\times[e]\subseteq\varphi$. Therefore, $\left\langle a,o,e\right\rangle\in\underline{\varphi}$. As a result, $\underline{X}\times\underline{Y}\times[e]\subseteq\underline{\varphi}$, and $e\in\underline{\alpha}(\underline{\varphi})_{E}(X,Y)$. ∎
Although they can be well-defined, the lower approximation operators seem to be less sensible that the upper approximation operators, and may sometimes lead to paradoxical conclusions.
Example 6.8.
We slightly change the setting of previous examples, yet still remain in the realm of pets.
  • We fix $A\coloneqq\text{the set of pets}\eqqcolon O$. Let $E$ be a set of environments.
  • Take the equivalence class: $\text{Koka}]$ (my former cat).
  • Let $X\coloneqq[\text{Koka}]\setminus\{\text{Koka}\}$ and $F\subseteq O$ be the set of fish in $O$.
  • Let $\varphi$ be an affordance of chasing.
Then, $\underline{\alpha}(\varphi)_{E}(X,F)=E$, which could be interpreted as: every cat similar to Koka (except for Koka herself) chases every fish from $F$ in every environment from $E$, due to the fact that $\underline{X}=\emptyset$, and so $\underline{X}\times\underline{F}\times\{e\}\subseteq\varphi$, for any $e\in E$. $\dashv$
The example above epitomizes a more general phenomenon. Let’s start with an equivalence class $K$. It may contain many elements about which we know much. But if we remove one $x$ from $K$, then $\underline{K}=\emptyset$, so in a way we lose all information about this class by approximating it from below. This may be an argument for focusing on upper approximations.
7. Mixing modal and approximations operators
@@
(7.3) Assume that $e\in[[\underline{\varphi}]]_{E}(X,Y)$, i.e., $X\times Y\times\{e\}\subseteq\underline{\varphi}$. We want to show that $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$. To this end, let $\left\langle x,y,f\right\rangle\in\underline{X}\times\underline{Y}\times[e]$. Thus $x]\subseteq X$, $y]\subseteq\underline{Y}$, and $[f]=[e]$. Therefore, $x\in X$ and $y\in Y$, and by the assumption we have that $\left\langle x,y,e\right\rangle\in\underline{\varphi}$, which means that $x]\times[y]\times[e]\subseteq\varphi$. In particular, $\left\langle x,y,f\right\rangle\in\varphi$, as desired.
@@
holds true. To show that it fails, we need to demonstrate that for some $X\subseteq A$, $Y\subseteq O$ and $e\in E$, $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$, yet $X\times Y\times\{e\}\nsubseteq\underline{\varphi}$, i.e., there are $x\in X$ and $y\in Y$ such that $x]\times[y]\times[e]\nsubseteq\varphi$. Again, let us use the ambient from Example 6.4. If we consider $r\coloneqq\mathrm{Robin}$ and its equivalence class $[r]$ (the set of all pets similar to him), then for $\coloneqq[r]\setminus\{r\}, X¯=\underline{X}$$\emptyset$. In consequence, we have that $\underline{\alpha}_{E}(\varphi)(X,Y)=E$. But, if we take $s$ to be a particular stick, and $e$ a particular dog park, and $\varphi$ an affordance of fetching, then it obviously does not have to be the case that $r]\times[s]\times[e]\subseteq\varphi$, i.e., there may be a dog similar to Robin and a stick similar to $s$ such that the dog cannot fetch the stick in $e$ (in our case, e.g., due to certain hidden factors not accessible from the point of vie of information systems). $\dashv$
@@
Similarly, for the sufficiency operator, we have that

### v2

@@ 5. Reasoning with affordances: modal-style operators
Having conceptualized affordances, we then turn to and focus on the problem of reasoning with them as ternary relations. As we will see, the equivalence classes of entities from three different sorts will play a crucial role in incorporating the granularity of perceptions of the world into the formal framework. In this section however, we begin with the definitions of crisp modal-style operators. As such these are independent from the affordance environment above, and can be laid out in a completely abstract setting. The logic we envisage is the binary version of an extension of classical propositional modal logic with a possibility operator and its dual necessity operator [1]. The expressive power of this logic is limited: even though it can express relational properties such as reflexivity, symmetry, and transitivity, it cannot express, for example, irreflexivity. To extend the expressive power of the classical logic, a “sufficiency” counterpart to the classical necessity operator was independently introduced by Humberstone, [15] and Gargov et al., [12]. The algebraic properties of such an operator and representation properties of such a “possibility-sufficiency logic” were investigated for the unary case by Düntsch et al., [11] and extended to the binary case by Düntsch et al., [10] where further details can be found. Besides the current investigation of affordances, a further example of an application of this logic is the study of formalizing a notion of betweeness [9]. We use the semantics of the logic developed in [10] to investigate properties of affordance relations. The setup we shall use is the following:
@@
The dual operators of $\left\langle\varphi\right\rangle$ and $\varphi]]$ are, respectively, $\varphi]_{E}$ and $\left\langle\left\langle\varphi\right\rangle\right\rangle_{E}$ , where
@@
Let us use one more time the interpretation from Example 5.1. Then,
@@
The second point confirms the awkwardness of the dual of the sufficiency operator. $\dashv$
Analogously to [9, Lemma 5.2] we have the conditions for the operators associated with respective sets.
@@
Let us look at the concrete example of a ternary relation in the form of betweenness that was studied by us in the setting of possibility-sufficiency algebras in [9]. $B(a,o,e)$ is interpreted as follows: the point $o$ is between (in the non-strict sense) the points $a$ and $e$. Then $e\in\left\langle B\right\rangle\rangle(X,Y)$ if and only if either there are $x\in X$ and $y\in Y$ such that $B(x,e,y)$ or either there are $x\in X$ and $y\in Y$ such that $\neg B(x,e,y)$. Then $\left\langle B\right\rangle\rangle(X,Y)=\emptyset$ if and only if at least one of the sets is empty, and $\left\langle B\right\rangle\rangle(X,Y)$ is the set of all points otherwise. $\dashv$
@@
1. Let $e\in E$. It is obvious that for any $x\in X$ and $o\in O$, if $\left\langle x,o,e\right\rangle\in\varphi$, then $x\in X$ or $o\in O$. From (5.7), $e\in[\varphi]_{E}(X,O)$, i.e. $\varphi]_{E}(X,O)=E$. In an analogous way we obtain that $\varphi]_{E}(A,Y)=E$.
@@ 6. The approximation operators
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(o)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\overline{\varphi}\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;([a]\times[o]\times[e])\cap\varphi\neq\emptyset.$
@@
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(o)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\underline{\varphi}\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;[a]\times[o]\times[e]\subseteq\varphi.$
@@
$\left\langle\left\langle a,f_{A}(a)\right\rangle,\left\langle o,f_{O}(o)\right\rangle,\left\langle e,f_{E}(e)\right\rangle\right\rangle\in\overline{\varphi}$
@@
A rough affordance is a pair $\left\langle\underline{\varphi},\overline{\varphi}\right\rangle$ where $\varphi$ is an affordance relation. $\dashv$
@@
With an affordance $\varphi$ we can associate its upper approximation $\overline{\alpha}_{S}(\varphi)$ on every coordinate $S\in\{A,O,E\}$ i.e., we have approximations for, respectively, actors, objects, and end environments:
@@
Let us comment on the relation of the $\overline{\alpha}$-operators to possibility operators from page 5. For this, we fix $S$ to be $E$. Both the possibility operator $\left\langle\varphi\right\rangle_{E}$ and the approximation operator $\overline{\alpha}_{E}(\varphi)$ have the type $2^{A}\times 2^{O}\to 2^{E}$, but they behave differently. The former is crisp and sends the pairs of collections of agents and objects, respectively, to the collection of environments in which agents and objects may $\varphi$-interact. The latter is rough, as given a pair $\left\langle X,Y\right\rangle$, where $X$ is a collection of agents and $Y$ is a collection of objects, $\overline{\alpha}_{E}(X,Y)$ is the collection of all those environments (contexts) $e$ such that an $X$-like agent and $Y$-like object $\varphi$-interact in an environment resembling $e$.
@@
(6.1) $\overline{\alpha}_{S}(\varphi)=\overline{\alpha}_{S}\left(\overline{\varphi}\right).$
@@
($\subseteq$) Let $e\in\overline{\alpha}_{E}(\varphi)(X,Y)$, i.e., $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\varphi\neq\emptyset$. Then, by Lemma 6.2 it must be the case that $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. So $e\in\overline{\alpha}_{E}(\overline{\varphi})(X,Y)$.
($\supseteq$) For the other direction, assume that $e\in\overline{\alpha}_{E}(\overline{\varphi})(X,Y)$, that is $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. Let $x\in\overline{X}$, $y\in\overline{Y}$, and $z\in[e]$ be such that $([x]\times[y]\times[z])\cap\varphi\neq\emptyset$. Let $\left\langle a,o,u\right\rangle$ be a triple which is in the intersection. But then $[a]=[x]$ and $[o]=[y]$ and $[u]=[e]$, and so $a\in\overline{X}$, $o\in\overline{Y}$, and in consequence $\left(\overline{X}\times\overline{Y}\times[e]\right)\cap\overline{\varphi}\neq\emptyset$. Thus $e\in\overline{\alpha}_{E}(\varphi)(X,Y)$, as desired. ∎
@@
  • $r\in\overline{\alpha}^{1}_{A}(\varphi)(S,P)$ means that a dog similar to Robin fetches a stick in a dog park, that is we are certain what the object and the environment involved in the action are, but we are unsure with respect to the actor.
  • $r\in\overline{\alpha}^{2}_{A}(\varphi)(S,P)$ means that Robin fetches a stick-like object in a dog park, so the actor and the environment are crisp, but the object is only approximated.
  • $r\in\overline{\alpha}^{3}_{A}(\varphi)(S,P)$ means that Robin fetches a stick in an environment resembling a dog park. Here the uncertainty concerns the environment, while we are unequivocal about the actor and the object.
  • $r\in\overline{\alpha}^{12}_{A}(\varphi)(S,P)$ means that in a dog park a dog similar to Robin fetches an object similar to a stick. Thus, the level of roughness rises as we only approximate two out of the three components of the action: the environment is crisp, but the actor and the object are rough.
  • $r\in\overline{\alpha}^{13}_{A}(\varphi)(S,P)$ means that in an environment resembling a dog park a dog similar to Robin fetches a stick. The situation is as above, with certainty shifted to the object, with the two remaining components rough.
  • $r\in\overline{\alpha}^{23}_{A}(\varphi)(S,P)$ means that Robin fetches something similar to a stick in an environment resembling a dog park. We are sure of the actor’s identity, but we can only surmise what the object and environment are.
Playing around with the coordinates, we can express the roughness of various aspects of the scenarios we put into focus. Clearly, approximating all three coordinates results in $\overline{\alpha}_{A}(\varphi)$ operator. $\dashv$
Example 6.6.
Imagine yourself going through the family album and encountering a photograph of your maternal grandparents, with your grandfather ($a$) kissing your grandmother ($b$) on the cheek. The background is a medieval town, which you know is in Italy, and which you find similar to Urbino. In this case, assuming that $\varphi$ is the affordance of kissing on the cheek, $\overline{\alpha}_{E}^{3}(\varphi)(\{a\},\{b\})$ is the collection of all medieval Italian towns that are similar to Urbino. $\dashv$
Example 6.7.
The manakins are small birds native to Central America distinguished by a dancing ritual involving a kind of moonwalk, usually performed on a tree branch [19]. An ornithologist fascinated by the phenomenon may naturally ask: are there similar birds mating in this way? To find out, they should localize places (environments) where a male shows off in an analogous way. To capture it in our framework, we set the set $A$ of actors, and $O$ of objects to consist of birds (so $A=O$). Let $M$ be the set of manakin males, $F$ of manakin females. Let $\varphi$ be the affordance of moonwalking in front of your potential partner, limited to birds. With this interpretation we have that
$\overline{\alpha}^{12}_{E}(M,F)=\{e\in E\mid(\overline{M}\times\overline{F}\times\{e\})\cap\varphi\neq\emptyset\}$
is the collection of all environments in which birds similar to manakins participate in the dancing ritual depicted in the linked footage, i.e., the places of interest of the curious ornithologist. $\dashv$
For any $i\in\{1,2,3,12,13,23\}$ by Lemma 6.2 we have
@@
However, unlike the unrestricted versions in (6.1), the reverse inclusions fail in general, as we show below.
@@
We stick to the interpretation from Example 6.4, and we consider the case in which $i\coloneqq 23$. Suppose that $r\in\overline{\alpha}^{23}_{A}(\overline{\varphi})(S,P)$, i.e.
$r\in\overline{\alpha}^{23}_{A}(\overline{\varphi})(S,P)$ $\Longleftrightarrow(\{r\}\times\overline{S}\times\overline{P})\cap\overline{\varphi}\neq\emptyset,$ by Def. of $\overline{\alpha}^{23}_{A}$,
$\Longleftrightarrow(\exists s\in S,p\in P)[(\{r\}\times[s]\times[p])\cap\overline{\varphi}\neq\emptyset],$ by Def. of $\overline{X}$,
$\Longleftrightarrow(\exists s\in S,p\in P)[([r]\times[s]\times[p])\cap\varphi\neq\emptyset],$ by Def. of $\overline{\varphi}$.
This tells us that some dog resembling Robin ($r^{\prime}\in[r]$) fetches a stick-like object ($o^{\prime}\in[o]$) in a dog-park-like environment ($p^{\prime}\in[p]$). If $|[r]|\geqslant 2$, then we cannot say for sure that the dog $r^{\prime}$ that resembles Robin is, indeed, Robin. To see this, observe that by the definition we have
$r\in\overline{\alpha}^{23}_{A}(\varphi)(S,P)\Longleftrightarrow(\{r\}\times\overline{S}\times\overline{P})\cap\varphi\neq\emptyset.$
Observe that in the equivalence above $\varphi$ is used and not its upper approximation, and that the first projection is fixed. Therefore, we may only take into account $\{r\}$, but in general not its equivalence class $[r]$. In the fleshy entourage of the example, we attribute certainty to the actor’s identity, i.e., Robin. Thus, passing from $r\in\overline{\alpha}^{23}_{A}(\overline{\varphi})(S,P)$ to $r\in\overline{\alpha}^{23}_{A}(\varphi)(S,P)$ expressed in natural language is like saying “From the fact that a dog similar to Robin is involved in an action, I can conclude that Robin himself is”. But, as we observed, it cannot be done if the collection of dogs similar to Robin has more than one element. $\dashv$
@@
For each $S\in\{A,O,E\}$ we have
$\underline{\alpha}_{S}(\varphi)(X,Y)=\underline{\alpha}_{S}(\underline{\varphi})_{S}(X,Y).$
@@
This time, the inclusion $\supseteq$ is immediate in light of Lemma 6.2. For $\subseteq$, assume that $e\in\underline{\alpha}_{E}(\varphi)(X,Y)$, i.e., $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$. Let $\left\langle a,o,f\right\rangle$ be from $\underline{X}\times\underline{Y}\times[e]$. Thus $a]\subseteq X$, $o]\subseteq Y$, and $[f]=[e]$. So, $a]\subseteq\underline{X}$, $o]\subseteq\underline{Y}$, and $a]\times[o]\times[e]\subseteq\varphi$. Therefore, $\left\langle a,o,e\right\rangle\in\underline{\varphi}$. As a result, $\underline{X}\times\underline{Y}\times[e]\subseteq\underline{\varphi}$, and $e\in\underline{\alpha}_{E}(\underline{\varphi})(X,Y)$. ∎
However, as we are mainly interested in upper approximations epitomizing roughness and imprecision phenomena, here, we do not delve into properties of the lower approximation operators leaving this topic for further work.
7. Mixing modal and approximation operators
@@
(7.3) Assume that $e\in[[\underline{\varphi}]]_{E}(X,Y)$, i.e., $X\times Y\times\{e\}\subseteq\underline{\varphi}$. We want to show that $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$. To this end, let $\left\langle x,y,f\right\rangle\in\underline{X}\times\underline{Y}\times[e]$. Thus $x]\subseteq X$, $y]\subseteq\underline{Y}$, and $[f]=[e]$. Therefore, $x\in X$ and $y\in Y$, and by the assumption we have $\left\langle x,y,e\right\rangle\in\underline{\varphi}$, which means that $x]\times[y]\times[e]\subseteq\varphi$. In particular, $\left\langle x,y,f\right\rangle\in\varphi$, as desired.
@@
holds true. To show that it fails, we need to demonstrate that for some $X\subseteq A$, $Y\subseteq O$ and $e\in E$, $\underline{X}\times\underline{Y}\times[e]\subseteq\varphi$, yet $X\times Y\times\{e\}\nsubseteq\underline{\varphi}$, i.e., there are $x\in X$ and $y\in Y$ such that $x]\times[y]\times[e]\nsubseteq\varphi$. Again, let us use the ambient from Example 6.4. If we consider $r\coloneqq\mathrm{Robin}$ and its equivalence class $[r]$ (the set of all pets similar to him), then for $\coloneqq[r]\setminus\{r\}, X¯=\underline{X}$$\emptyset$. In consequence, we have $\underline{\alpha}_{E}(\varphi)(X,Y)=E$. But, if we take $s$ to be a particular stick, and $e$ a particular dog park, and $\varphi$ an affordance of fetching, then it obviously does not have to be the case that $r]\times[s]\times[e]\subseteq\varphi$, i.e., there may be a dog similar to Robin and a stick similar to $s$ such that the dog cannot fetch the stick in $e$ (in our case, e.g., due to certain hidden factors not accessible from the point of vie of information systems). $\dashv$
@@
Similarly, for the sufficiency operator, we have